[The following information applies to the questions displayed below.]

In 2021, the Westgate Construction Company entered into a contract to construct a road for Santa Clara County for $10,000,000. The road was completed in 2023. Information related to the contract is as follows:

2021 2022 2023 Cost incurred during the year $ 2,580,000 $ 4,042,000 $ 2,175,800 Estimated costs to complete as of year-end 6,020,000 1,978,000 0 Billings during the year 2,060,000 4,562,000 3,378,000 Cash collections during the year 1,830,000 4,200,000 3,970,000

Westgate recognizes revenue over time according to percentage of completion.

5. Calculate the amount of revenue and gross profit (loss) to be recognized in each of the three years assuming the following costs incurred and costs to complete information. (Do not round intermediate calculations and round your final answers to the nearest whole dollar amount. Loss amounts should be indicated with a minus sign.) 2021 2022 2023

Costs incurred during the year $ 2,580,000 $ 3,830,000 $ 3,990,000 Estimated costs to complete as of year-end 6,020,000 4,160,000 0
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Westgate Construction Company
Calculation of Gross Profit
Particulars 2021 2022 2023
Contract Price         10,000,000       10,000,000         10,000,000 A
Cost Incurred during the year           2,580,000          4,042,000           2,175,800 B
Cost Incurred to date           2,580,000          6,622,000           8,797,800 C
(Cost incurred to date of the present year less Cost incurred to date of the previous year)
Estimated cost to complete           6,020,000          1,978,000                         -   D
Estimated total cost           8,600,000          8,600,000           8,797,800 E=C+D
(Cost Incurred to date+ Estimated cost to complete )
Percentage complete 30.00% 77.00% 100.00% F=C/E
(Cost Incurred to date/Estimated total cost)
Revenue to be recognized           3,000,000          4,700,000           2,300,000 G
(Contract Price*Percentage complete)- revenue previously recognized
Year 2021: 30% completed. Revenue recognized = 30% x $ 10,000,000 = $ 3,000,000.
Year 2022: 77% completed. Revenue recognized = 77% x $ 10,000,000 – $ 3,000,000 (previously recognized) = $ 4,700,000.
Year 2023: 100% completed. Revenue recognized = 100% x $ 10,000,000 – $ 3,000,000- $ 4,770,000 (previously recognized) = $ 2,300,000.
Gross Profit/(Loss)              420,000             658,000              124,200 H=G-B
Billings during the year     2,060,000.00    4,562,000.00     3,378,000.00
Cash collections during the year     1,830,000.00    4,200,000.00     3,970,000.00
Westgate Construction Company
Answer 5 Calculation of Gross Profit
Particulars 2021 2022 2023
Contract Price         10,000,000       10,000,000         10,000,000 A
Cost Incurred during the year           2,580,000          3,830,000           3,990,000 B
Cost Incurred to date           2,580,000          6,410,000         10,400,000 C
(Cost incurred to date of the present year less Cost incurred to date of the previous year)
Estimated cost to complete           6,020,000          1,978,000                         -   D
Estimated total cost           8,600,000          8,388,000         10,400,000 E=C+D
(Cost Incurred to date+ Estimated cost to complete )
Percentage complete 30.00% 76.42% 100.00% F=C/E
(Cost Incurred to date/Estimated total cost)
Revenue to be recognized           3,000,000          4,641,869           2,358,131 G
(Contract Price*Percentage complete)- revenue previously recognized
Year 2021: 30% completed. Revenue recognized = 30% x $ 10,000,000 = $ 3,000,000.
Year 2022: 76% completed. Revenue recognized = 76% x $ 10,000,000 – $ 3,000,000 (previously recognized) = $ 4,641,869.
Year 2023: 100% completed. Revenue recognized = 100% x $ 10,000,000 – $ 3,000,000- $ 4,641,869 (previously recognized) = $ 2,3581,131.
Gross Profit/(Loss)              420,000             811,869         (1,631,869) H=G-B